# Spontaneous emission and lasing in photonic time crystals

### v1

@@ Abstract
We report the first direct mapping of the local density of states (LDOS) in a photonic time crystal (PTC), capturing its evolution from the analogues of spontaneous emission enhancement to thresholded lasing. The PTC is implemented with an array of time‑periodically modulated LC resonators at microwave frequencies. Broadband white noise probes the system and reveals an LDOS lineshape that decomposes into absorptive and dispersive Lorentzian components near the momentum gap frequency. The finite peak amplitude, which grows with modulation strength, shows that the spontaneous emission rate is maximized at the gap frequency. All observed features agree with classical non‑Hermitian Floquet theory. When modulation‑induced gain exceeds losses, the PTC transitions to a narrow‑band lasing oscillation state. These findings open a route to nonequilibrium photonics and bring time‑periodic LDOS engineering closer to practical realization.
@@
Controlling spontaneous emission decay by reshaping the electromagnetic vacuum has guided photonics since Purcell’s pioneering cavity proposal [1], through photonic spatial crystals [2, 3, 4, 5, 6, 7, 8, 9] and, more recently, hyperbolic metamaterials [10, 11]. In every case the decisive parameter is the local photonic density of states (LDOS), which quantifies the vacuum field fluctuations available to a quantum emitter at a given frequency, position, and dipole orientation [12, 13, 14, 15]. Although static nanostructures have proven effective in shaping the LDOS in both space and frequency, their modal spectrum remains fixed and unchanging once fabricated.
@@
Photonic time crystals (PTCs) break this constraint by modulating the optical parameters such as refractive index or permittivity in time rather than in space [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Note that our use of the term time crystal differs from the conventional notion of a discrete time crystal, which refers to a many‑body quantum or classical phase that spontaneously breaks discrete time‑translation symmetry. A periodic modulation at frequency $\Omega$ folds the dispersion into Floquet replicas, mixes positive- and negative-frequency branches, and opens momentum gaps [31, 32, 33]. Classical electrodynamics predicts that, near the edges of a momentum gap, the Petermann factor enhances the modal contributions to the momentum‑resolved density of states (kDOS). Interestingly, in the presence of time‑periodicity-induced gain, the kDOS can even become negative inside the gap, foreshadowing unconventional effects such as spontaneous emission excitation [34]. These predictions are based on a classical non-Hermitian Maxwell-Floquet framework (which can also be inferred from full quantum electrodynamics [35]) that has no static analogue, though direct experimental validation has been lacking.
@@
In this Letter, we report the first direct measurement of the spectrally resolved LDOS and, by extension, the spontaneous emission spectrum, in a PTC. Operating in the linear but lossy regime, we measure the radiated power generated by broadband white noise in the driven circuits that constitute the PTC. The resulting spectrum exhibits a pronounced cusp at the momentum gap frequencies and can be decomposed into absorptive and dispersive Lorentzian components. All features agree with classical non-Hermitian Floquet theory, which predicts an enhancement of the LDOS, and therefore of the spontaneous emission decay rate, at the gap frequency [34]. These measurements provide the first direct evidence that purely temporal modulation can sculpt the electromagnetic environment of a light emitter. Furthermore, we observe a transition into a narrow-band, oscillating regime above a threshold modulation power—a regime we refer to as a PTC laser.
@@
Figure 1: (a) Schematic of the equivalent circuit for each unit cell in the array, featuring two degrees of freedom that facilitate the generation of two distinct LC resonances. In each unit cell, one of the three capacitors undergoes sinusoidal modulation at an angular frequency of $\Omega$, altering the effective permittivity of the time crystal in a time-periodic fashion. (b) Maps of momentum-resolved density of states along with band structures (main plots), LDOS (left panels) and imaginary quasi-eigenfrequencies (bottom panels) for $\Omega\approx 2\omega_{k}$ at $\delta=0.023$.
@@
We first analyze the band structure of the PTC emulated with a one‑dimensional array of LC resonators side‑coupled to a bus waveguide (Fig. 1a). The capacitance of each resonator is modulated periodically in time as $C_{r}(t)=C_{c}[1+\delta(t)]^{-1}$, where $\delta(t)$ is a dimensionless function with period $T=2\pi/\Omega$. In momentum space, the complex amplitudes of the waveguide and resonator modes, $a_{k}$ and $b_{k}$, evolve under a time‑dependent Hamiltonian $\mathcal{H}(t)=\mathcal{H}^{(0)}+\mathcal{H}^{(1)}(t)$. The static part $\mathcal{H}^{(0)}$ contains the bare dispersions of the waveguide and resonators and their mutual coupling, whereas the time-dependent part $\mathcal{H}^{(1)}(t)$ originates from the periodic capacitance drive. Defining the mode vector as $\mathbf{x}_{k}=[a_{k},b_{k},a_{-k}^{\dagger},b_{-k}^{\dagger}]^{T}$, the equations of motion reduce to the compact form $i\dot{\mathbf{x}}_{k}=\mathbf{H}_{k}(t)\mathbf{x}_{k}$, with $\mathbf{H}_{k}(t)=\mathbf{H}_{k}^{(0)}+\mathbf{H}_{k}^{(1)}(t)+\mathbf{D}_{k}$. To model the lossy regime, each resonator is equipped with a parallel resistor, which leads to a loss-induced term $\mathbf{D}_{k}$ in the dynamical matrix. Full expressions are provided in Supplementary Material B.
@@
For simplicity, we assume a single-harmonic modulation $\delta(t)=\delta\cos(\Omega t)$, where $\Omega$ is the modulation frequency. Under this assumption the Floquet Hamiltonian $\mathbf{H}_{k}^{F}$ becomes block-tridiagonal. The corresponding eigenvalue equation $\mathbf{H}_{k}^{F}\mathbf{F}_{k,\alpha}=\omega_{\alpha}\mathbf{F}_{k,\alpha}$, where $\mathbf{F}_{k,\alpha}$ denotes the Floquet eigenvector, defines the quasi-frequency band structure of the PTC. At the near-resonant condition $\Omega\simeq 2\omega_{k}$, with $\omega_{k}$ representing the Bloch‑mode frequency of the resonator array, asymmetric coupling occurs between the original positive-frequency band and the Floquet sideband of its corresponding negative-frequency band, opening a momentum gap at $\Omega/2$ (Fig. 1b). The plot combines continuum-limit Bloch–Floquet dispersions with discrete $k$-resolved resonances whose spacing reflects the finite number of unit cells ($N=12$). At the edges of these gaps, we observe exceptional points (EPs), where Floquet eigenmode and their quasi-eigenfrequencies coalesce simultaneously. Within the momentum gap, the two modes exhibit unequal decay rates, as reflected by the splitting in the imaginary parts of the quasi-eigenfrequencies (see the bottom panel of Fig. 1b).
Figure 2: Real (solid) and imaginary (dashed) parts of the normalized field intensity $I_{\pm}(k)$. Inside the momentum gap (green shading) the finite $\Im[I_{\pm}(k)]$ yields the anti-symmetric Lorentzian weight observed in e–g; outside the gap $\Im[I_{\pm}(k)]\approx 0$, leaving a purely symmetric line shape.
Figure 3: (a) Experimentally measured radiated-power map at the momentum gap frequency as functions of modulation frequency and power. Blue (linear) and yellow (oscillation/lasing) regions are separated by the oscillation threshold; colored markers indicate operating points used in the panels b and g. (b) Sub-threshold spectra measured for the gap-center modes. Increasing the modulation power from -17 dBm to -16 dBm increases the peak amplitude and narrows the symmetric Lorentzian profile. (c) The calculated LDOS for modulation depths in the range $0.023\leq\delta\leq 0.027$ reproduces the observed peak growth and linewidth narrowing. (d) Imaginary part of the eigenfrequencies as a function of wavenumber $k$. The two gap modes dominates the LDOS at the momentum gap frequency. (e) Real and imaginary parts of $I_{\alpha}$ for the gap modes as a function of modulation frequency. The imaginary parts of $I_{\alpha}$ decreases as the mode approaches the center of the momentum gap, reducing the anti-symmetric weight in the LDOS. (f) Extracted linewidth $\Gamma_{-}=\Im[\omega_{-}]$ of the lower-loss gap mode. The linewidth decreases toward the gap center, accompanied by a narrowing of the LDOS peak. (g) When the modulation frequency is detuned so the modes move off-center in the momentum gap, the measured spectrum becomes skewed. Black curve: total fit obtained from the two-mode model of Eq (3). Blue and orange curves: individual contributions of the $\alpha=+$ and -- gap modes, respectively. (h) Theoretical LDOS for this detuned case reproduces the asymmetric profile. (i) Corresponding Imaginary part of the eigenfrequencies versus wavenumber $k$.
@@
In the Floquet picture, the response of the system is governed by the Floquet Green function $G_{F}(k,\omega)=[\omega\mathbf{I}_{F}-\mathbf{H}_{k}^{F}]^{-1}$, where $\mathbf{H}_{k}^{F}$ contains all harmonic blocks ($q=0,\,\pm 1,\,\pm 2,\,\dots$) and loss. A monochromatic drive at frequency $\omega$ couples to the resonator mode $b_{k}$ through $\mathbf{s}_{0}=[0,-1,0,1]^{T}$ in the $q=0$ block, thereby exciting all side-band components $b_{k}^{(q)}$ at $\omega+q\Omega$. Embedding $\mathbf{s}_{0}$ in the full Floquet space, $\mathbf{s}=[\dots;\mathbf{0};\mathbf{s}_{0};\mathbf{0};\dots]$, we project the Floquet Green function onto the same ($b_{k},q=0$) channel:
@@
which is the Green function for the resonator mode $b_{k}$ at frequency $\omega$. Here, $g_{F}$ describes the flux response, dividing by the $k$-dependent effective inducance $L^{\prime}_{r,k}$ (see Supplementary Material B) converts it into the complex current-to-voltage response $Y_{F}(k,\omega)=g_{F}(k,\omega)/L^{\prime}_{r,k}$, i.e. the admittance at $(k,\omega)$. The poles $\omega_{\alpha}=\Omega_{\alpha}-i\Gamma_{\alpha}$ reproduce the quasi-eigenfrequencies, while $\Im[Y_{F}]$ gives the spectral weight that can absorb power from the drive. We therefore define the momentum-resolved density of states (kDOS)
@@
and obtain the LDOS by summation $\rho(\omega)=N^{-1}\sum_{k}\rho_{k}(\omega)$. Neglecting edge effects under the bulk approximation, we treat the LDOS as translationally invariant and omit the site index. The LDOS shows a sharp peak at the momentum gap frequency(Fig. 1b), confirming our recent prediction that spontaneous emission decay rates are enhanced at this spectral point[34]. If we consider only the two modes $\omega_{\pm}=\Omega_{\pm}-i\Gamma_{\pm}$ forming the momentum gap, the kDOS can be approximately decomposed as
@@
with $\mathbf{F}_{k,\alpha}^{R}$ and $\mathbf{F}_{k,\alpha}^{L}$ denoting the right and left eigenvectors of the Floquet Hamiltonian $\mathbf{H}_{k}^{F}$, respectively. Since the denominator involves the overlap between left and right eigenvectors, $I_{\alpha}(k)$ grows with the square root of the Petermann factor, capturing the effect of nonorthogonal eigenmodes. The real and imaginary parts of $I_{\alpha}(k)$ determine the weights of the symmetric and anti-symmetric Lorentzian components that constitute the kDOS. For momenta $k$ within the band region (i.e., outside the momentum gap), $\Im[I_{\alpha}(k)]$ is negligible, yielding a symmetric Lorentzian profile. In contrast, inside the momentum gap, $\Im[I_{\alpha}(k)]$ becomes significant, introducing an anti-symmetric contribution to the kDOS (Fig. 2). The weight of the anti-symmetric component increases as the modes approach the gap edges and vanishes at the gap center.
@@
To implement the unit cell of the PTC, a metallic split ring resonator (SRR) was patterned onto a printed circuit board. Within this unit cell, we incorporated a varactor— an element whose capacitance changes when the voltage across it is varied. By applying a DC-biased AC voltage to the varactor, the resonant characteristic of the unit cell can be time-periodically modulated. To build a PTC, we inserted an array of 12 unit cells into the waveguide. To verify the predicted LDOS enhancement at the momentum gap frequency, we excite the PTC with broadband white noise arising from Johnson–Nyquist thermal fluctuations along with residual technical noise. These fluctuations plays the role of a bath of classical emitters distributed throughout the PTC. The measured radiated-power spectrum, which is proportional to the spontaneous emission decay rate, matches, with high fidelity, the LDOS profile obtained from our non-Hermitian Floquet theory.
@@
Figure 3a shows an experimentally measured map of the oscillation power at the momentum gap frequency as a function of the modulation power and frequency. Once the modulation power exceeds the intrinsic loss in the PTC, varactor nonlinearity drives the system into a saturated oscillating (lasing) regime that lies beyond the scope of linear Floquet analysis, as discussed later. We begin by analyzing the linear regime, where the Floquet description remains valid. To ensure operation within this linear, lossy response regime, we systematically reduced the modulation power below the oscillation threshold, recorded the radiated power spectral density or correspondingly the spectrally resolved LDOS (Fig. 3b, g).
@@
Figure 4: (a) Measured radiated power at the momentum gap frequency and linewidth as a function of modulation power at fixed modulation frequency $\Omega/2\pi=2.998\,\mathrm{GHz}$. As the modulation power exceeds the threshold, the system undergoes a sharp transition into an oscillating regime where the radiated power saturates due to nonlinearities. (b) Spectral profile at $-13\,\mathrm{dBm}$, deep in the lasing regime, showing a sharp, narrow-band oscillation.
@@
In the experiment, the position of the modes within the momentum gap can be tuned by adjusting the modulation frequency $\Omega$. We begin with the case where $\Omega$ is set to place the modes at the center of the gap. With this frequency fixed, we measure the spectrum while sweeping the modulation power by adjusting the peak-to-peak AC voltage applied to the varactor diodes, each biased at a fixed reverse DC voltage. The measured spectra in Fig.3b are well fitted by a symmetric Lorentzian function, with increasing modulation power leading to higher peak amplitude and narrower linewidth. This behavior is closely reproduced by the theoretical curves in Fig.3c,d. When the modes are displaced from the center of the momentum gap, the measured LDOS becomes noticeably asymmetric with respect to the gap frequency. This skewness is fully consistent with the our prediction that the spectrum consists of symmetric and anti-symmetric Lorentzian components (Fig. 3g-i).

### v2

# Analogs of spontaneous emission and lasing in photonic time crystals

@@ Abstract
We report the first direct mapping of the frequency‑resolved local density of states (LDOS) in a photonic time crystal (PTC) implemented as an array of time‑periodically modulated LC resonators at microwave frequencies. Broadband white noise probes the system and yields an LDOS lineshape near the momentum gap that can be decomposed into absorptive and dispersive Lorentzian components. The finite LDOS peak at the gap frequency, which grows with modulation strength, implies that the spontaneous emission rate of an emitter coupled to the PTC would be maximized at that frequency. The measured spectra are in good agreement with classical non‑Hermitian Floquet theory. As the modulation‑induced gain exceeds intrinsic losses, the system undergoes a transition to a narrow-band self-oscillation (lasing) regime. These results open a route to nonequilibrium photonics and bring time‑periodic LDOS engineering closer to practical realization.
@@
Controlling spontaneous emission rates by reshaping the electromagnetic vacuum has been a central theme in nanophotonics since Purcell’s pioneering cavity proposal [PhysRev.69.674.2], and has been pursued in photonic crystals [1972JETP...35..269B, PhysRevLett.58.2059, john1987strong, fan1997high, PhysRevLett.95.013904, PhysRevLett.99.023902, PhysRevLett.99.193901, Noda2007] and, more recently, hyperbolic metamaterials [jacob2012broadband, noginov2010controlling]. In these platforms, within the weak‑coupling regime, the spontaneous emission rate of a dipole emitter is governed by the local photonic density of states (LDOS), which quantifies the vacuum field fluctuations available to a quantum emitter at a given frequency, position, and dipole orientation [lodahl2004controlling, vos2009orientation, novotny2012principles, milonni2013quantum]. Although static, passive nanostructures can shape the LDOS in both space and frequency, their modal spectrum is fixed by geometry and material parameters once fabricated.
@@
Photonic time crystals (PTCs) break this constraint by modulating the optical parameters such as refractive index or permittivity in time rather than relying on spatial structuring [zurita2009reflection, salem2015temporal, PhysRevA.93.063813, doi:10.1063/1.4928659, PhysRevB.98.085142, martinez2018parametric, Chamanara2018, Park2021, Lee2021, asgari2024theory_, wang2024expanding, sustaeta2025quantum, dong2025extremely]. Note that our use of the term time crystal differs from the conventional notion of a discrete time crystal, which refers to a many‑body quantum or classical phase that spontaneously breaks discrete time‑translation symmetry. A periodic modulation at frequency $\Omega$ folds the dispersion into Floquet replicas that hybridize positive‑ and negative‑frequency branches and open momentum gaps [doi:10.1126/sciadv.abo6220, doi:10.1126/sciadv.adg7541, feinberg2025plasmonic]. Classical electrodynamics predicts that, near the edges of a momentum gap, the Petermann factor, which quantifies Floquet-mode non-orthogonality, enhances the contributions of individual modes to the momentum-resolved density of states (kDOS). Interestingly, in parameter regimes with net time-periodicity-induced gain, the kDOS can become negative, foreshadowing unconventional effects such as spontaneous emission excitation [park2025spontaneous]. These predictions are based on a classical non-Hermitian Maxwell–Floquet framework and can also be inferred from a quantum electrodynamical description [bae2025cavity], although the corresponding kDOS and LDOS have not yet been directly tested experimentally.
@@
In this Letter, we report what is, to our knowledge, the first direct measurement of the spectrally resolved LDOS and, by extension, the spontaneous‑emission spectrum of an emitter coupled to a PTC. Operating in the linear but lossy regime, we measure the radiated power generated by broadband white noise in the driven circuits that constitute the PTC. The resulting spectrum exhibits a pronounced cusp at the momentum gap frequency and can be decomposed into absorptive and dispersive Lorentzian components. All features agree well with classical non‑Hermitian Floquet theory, which predicts an enhancement of the LDOS, and therefore of the spontaneous‑emission decay rate, at the gap frequency [park2025spontaneous]. These measurements provide the first direct evidence that purely temporal modulation can sculpt the electromagnetic environment of a light emitter. Furthermore, above a threshold modulation power we observe a transition to a narrow-band oscillation consistent with modulation-induced parametric self-oscillation (hereafter a PTC laser).
@@
Figure 1: (a) Schematic representation of the equivalent circuit for the photonic time crystal. The system consists of an array of LC resonators side‑coupled to a transmission line. In each unit cell, the resonator capacitors are sinusoidally modulated at angular frequency $\Omega$, thereby producing a time‑periodic modulation of the effective permittivity of the structure. (b) Map of the momentum-resolved density of states (kDOS) together with band structures (main panel), LDOS (left panel), and imaginary parts of the quasi-eigenfrequencies (bottom panel) for $\Omega\approx 2\omega_{k}$ at $\delta=0.023$.
@@
We first analyze the band structure of the PTC implemented as a one‑dimensional array of LC resonators side‑coupled to a bus waveguide (Fig. 1a). The capacitance of each resonator is modulated periodically in time as $C_{r}(t)=C_{c}[1+\delta(t)]^{-1}$, where $\delta(t)$ is a dimensionless function with period $T=2\pi/\Omega$. In momentum space, the complex amplitudes of the waveguide and resonator modes, $a_{k}$ and $b_{k}$, evolve under a time‑dependent Hamiltonian $\mathcal{H}(t)=\mathcal{H}^{(0)}+\mathcal{H}^{(1)}(t)$. The static part $\mathcal{H}^{(0)}$ contains the bare dispersions of the waveguide and resonators and their mutual coupling, whereas the time-dependent part $\mathcal{H}^{(1)}(t)$ originates from the periodic capacitance drive. Defining the mode vector as $\mathbf{x}_{k}=[a_{k},b_{k},a_{-k}^{\dagger},b_{-k}^{\dagger}]^{T}$, the equations of motion reduce to the compact form $i\dot{\mathbf{x}}_{k}=\mathbf{H}_{k}(t)\mathbf{x}_{k}$, with $\mathbf{H}_{k}(t)=\mathbf{H}_{k}^{(0)}+\mathbf{H}_{k}^{(1)}(t)+\mathbf{D}_{k}$. To model the lossy regime, each resonator is equipped with a resistor, which leads to a loss-induced term $\mathbf{D}_{k}$ in the dynamical matrix. Full expressions are provided in Supplementary Material B.
@@
For simplicity, we assume a single-harmonic modulation $\delta(t)=\delta\cos(\Omega t)$, where $\Omega$ is the modulation frequency. Under this assumption, the Floquet Hamiltonian $\mathbf{H}_{k}^{F}$ is block-tridiagonal. The corresponding eigenvalue equation $\mathbf{H}_{k}^{F}\mathbf{F}_{k,\alpha}=\omega_{\alpha}\mathbf{F}_{k,\alpha}$, where $\mathbf{F}_{k,\alpha}$ denotes the Floquet eigenvector, defines the quasi-frequency band structure of the PTC. At the near-resonant condition $\Omega\simeq 2\omega_{k}$, with $\omega_{k}$ representing the Bloch‑mode frequency of the resonator array, asymmetric coupling occurs between the original positive-frequency band and the Floquet sideband of its corresponding negative-frequency band, opening a momentum gap around $\Omega/2$ (Fig. 1b). The plot combines continuum-limit Bloch–Floquet dispersions with discrete $k$-resolved resonances whose spacing reflects the finite number of unit cells ($N=12$). At the edges of this gap, we observe exceptional points (EPs), where Floquet eigenmodes and their complex quasi-eigenfrequencies coalesce. Within the momentum gap, the two modes exhibit unequal decay rates, as reflected by the splitting in the imaginary parts of the quasi-eigenfrequencies (see the bottom panel of Fig. 1b).
Figure 2: (a) Real (solid lines) and imaginary (dashed lines) parts of the normalized field intensity $I_{\pm}(k)$. Inside the momentum gap (green shaded region), a finite $\Im[I_{\pm}(k)]$ generates the antisymmetric Lorentzian component in the kDOS. Outside the gap, $\Im[I_{\pm}(k)]\approx 0$, resulting in a purely symmetric Lorentzian profile. (b) Square root of the Petermann factor, $\sqrt{\mathrm{PF}}=\sqrt{\mathrm{PF}_{+}}=\sqrt{\mathrm{PF_{-}}}$ (black), and source-mode overlap magnitude $\Pi_{\alpha}$ for the two modes (++ in blue and - in red)
@@
In the Floquet picture, the response of the system is governed by the Floquet Green’s function $G_{F}(k,\omega)=[\omega\mathbf{I}_{F}-\mathbf{H}_{k}^{F}]^{-1}$, where $\mathbf{H}_{k}^{F}$ contains all harmonic blocks ($q=0,\,\pm 1,\,\pm 2,\,\dots$) and loss. A monochromatic drive at frequency $\omega$ couples to the resonator mode $b_{k}$ through $\mathbf{s}_{0}=[0,-1,0,1]^{T}$ in the $q=0$ block, thereby exciting all sideband components $b_{k}^{(q)}$ at $\omega+q\Omega$. Embedding $\mathbf{s}_{0}$ in the full Floquet space, $\mathbf{s}=[\dots;\mathbf{0};\mathbf{s}_{0};\mathbf{0};\dots]$, we project the Floquet Green’s function onto the same ($b_{k},q=0$) channel:
@@
which is the Green’s function for the resonator mode $b_{k}$ at frequency $\omega$. Here, $g_{F}$ describes the flux response; dividing by the $k$-dependent effective inductance $L^{\prime}_{r,k}$ (see Supplementary Material B) converts it into the complex current-to-voltage response $Y_{F}(k,\omega)=g_{F}(k,\omega)/L^{\prime}_{r,k}$, i.e., the admittance at $(k,\omega)$. The poles $\omega_{\alpha}=\Omega_{\alpha}-i\Gamma_{\alpha}$ reproduce the quasi-eigenfrequencies, while $\Im[Y_{F}]$ gives the spectral weight that can absorb power from the drive. We therefore define the momentum-resolved density of states (kDOS)
@@
and obtain the LDOS by summation $\rho(\omega)=N^{-1}\sum_{k}\rho_{k}(\omega)$. Neglecting edge effects under the bulk approximation, we treat the LDOS as translationally invariant and omit the site index. The LDOS shows a sharp peak at the momentum gap frequency (Fig. 1b), confirming our recent prediction that spontaneous emission decay rates are enhanced at this spectral point [park2025spontaneous]. If we consider only the two modes $\omega_{\pm}=\Omega_{\pm}-i\Gamma_{\pm}$ forming the momentum gap, the kDOS can be approximately decomposed as
@@
with $\mathbf{F}_{k,\alpha}^{R}$ and $\mathbf{F}_{k,\alpha}^{L}$ denoting the right and left eigenvectors of the Floquet Hamiltonian $\mathbf{H}_{k}^{F}$, respectively. Since the denominator involves the overlap between left and right eigenvectors, the magnitude $|I_{\alpha}(k)|$ grows with the square root of the Petermann factor, capturing the effect of non-orthogonal eigenmodes. The real and imaginary parts of $I_{\alpha}(k)$ determine the weights of the symmetric and antisymmetric Lorentzian components that constitute the kDOS. For momenta $k$ within the band region (i.e., outside the momentum gap), $\Im[I_{\alpha}(k)]$ is negligible, yielding a symmetric Lorentzian profile. In contrast, inside the momentum gap, $\Im[I_{\alpha}(k)]$ becomes significant, introducing an antisymmetric contribution to the kDOS (Fig. 2a). The weight of the antisymmetric component increases as the modes approach the gap edges and vanishes at the gap center.
To further clarify the contribution of non-orthogonality to the modal weights, Fig. 2b plots the square root of the Petermann factor $\sqrt{\mathrm{PF}_{\alpha}}$ and the source–mode overlap $\Pi_{\alpha}$. As shown, the overlap magnitude $\Pi_{\alpha}\equiv|\mathbf{s}^{T}\mathbf{F}_{k,\alpha}^{R}\mathbf{F}_{k,\alpha}^{L}\mathbf{s}|$ remains nearly constant throughout the momentum gap (shaded region), whereas $\sqrt{\mathrm{PF}_{\alpha}}$ sharply increases toward the gap edges. Because the normalized intensity satisfies
$|I_{\alpha}|=2\sqrt{\mathrm{PF}_{\alpha}}\,\Pi_{\alpha},$ (5)
the variation of $|I_{\alpha}(k)|$ inside the gap (Fig. 2a) directly reflects the increase of $\sqrt{\mathrm{PF}_{\alpha}}$. Thus, the enhancement of the antisymmetric contribution to the kDOS originates from the growing non-orthogonality of the two in-gap modes near the exceptional points at the edges of the momentum gap.
@@
To implement the unit cell of the PTC, a metallic split-ring resonator (SRR) was patterned onto a printed circuit board. Within this unit cell, we incorporated a varactor—an element whose capacitance changes when the voltage across it is varied. By applying a DC-biased AC voltage to the varactor, the resonant characteristic of the unit cell can be time-periodically modulated. To build a PTC, we inserted an array of 12 unit cells into the waveguide. Further details of the fabricated LC resonator and the experimental setup are provided in Supplementary Material D. To verify the predicted LDOS enhancement at the momentum-gap frequency, we operate the PTC with no coherent input; the only drive is broadband white noise arising from Johnson–Nyquist thermal fluctuations together with residual technical noise. These fluctuations act as spatially distributed, broadband noise sources that probe the linear response of the PTC. The measured noise-driven radiated spectrum directly follows the LDOS lineshape predicted by non-Hermitian Floquet theory, providing an experimental proxy for the frequency dependence of the spontaneous-emission rate in the weak-coupling limit.
Figure 3: (a) Experimentally measured radiated-power map at the momentum gap frequency as functions of modulation frequency and power. Blue (linear) and yellow (oscillation/lasing) regions are separated by the oscillation threshold; colored markers indicate operating points used in the panels b and g. (b) Sub-threshold spectra measured for the gap-center modes. Increasing the modulation power from -17 dBm to -16 dBm increases the peak amplitude and narrows the symmetric Lorentzian profile. (c) The calculated LDOS for modulation depths in the range $0.023\leq\delta\leq 0.027$ reproduces the observed peak growth and linewidth narrowing. (d) The imaginary part of the eigenfrequencies as a function of wavenumber $k$. The two gap modes dominate the LDOS at the momentum gap frequency. (e) Real and imaginary parts of $I_{\alpha}$ for the gap modes as a function of modulation frequency. The imaginary parts of $I_{\alpha}$ decrease as the mode approaches the center of the momentum gap, reducing the antisymmetric weight in the LDOS. (f) Extracted linewidth $\Gamma_{-}=\Im[\omega_{-}]$ of the lower-loss gap mode. The linewidth decreases toward the gap center, accompanied by a narrowing of the LDOS peak. (g) When the modulation frequency is detuned so the modes move off-center in the momentum gap, the measured spectrum becomes skewed. Black curve: total fit obtained from the two-mode model of Eq. (3). Blue and orange curves: individual contributions of the $\alpha=+$ and - gap modes, respectively. (h) Theoretical LDOS for this detuned case reproduces the asymmetric profile. (i) Corresponding imaginary part of the eigenfrequencies versus wavenumber $k$.
@@
Figure 3a shows an experimentally measured map of the oscillation power at the momentum gap frequency as a function of the modulation power and frequency. Once the modulation crosses threshold, the array exhibits self-oscillation; in the above-threshold regime the varactor nonlinearity becomes essential and leads to saturation, beyond linear Floquet theory. We begin by analyzing the linear regime, where the Floquet description remains valid. To ensure operation within this linear, lossy response regime, we systematically reduced the modulation power below the oscillation threshold and recorded the radiated power spectral density, which corresponds to the spectrally resolved LDOS (Fig. 3b, g).
@@
Figure 4: (a) Measured radiated power at the momentum gap frequency and the corresponding linewidth as functions of modulation power at a fixed modulation frequency $\Omega/2\pi=2.998\,\mathrm{GHz}$. As the modulation power exceeds the threshold, the system undergoes a sharp transition into an oscillating regime where the radiated power saturates due to nonlinearities. (b) Spectral profile at $13\,\mathrm{dBm}$, deep in the lasing regime, showing a sharp, narrow-band oscillation.
@@
In the experiment, the position of the modes within the momentum gap can be tuned by adjusting the modulation frequency $\Omega$. We begin with the case where $\Omega$ is set so that the modes lie at the center of the gap. With this frequency fixed, we measure the spectrum while sweeping the modulation power by adjusting the peak-to-peak AC voltage applied to the varactor diodes, each biased at a fixed reverse DC voltage. The measured spectra in Fig. 3b are well fitted by a symmetric Lorentzian function, with increasing modulation power leading to a higher peak amplitude and a narrower linewidth. This behavior is closely reproduced by the theoretical curves in Figs. 3c,d. When the modes are displaced from the center of the momentum gap, the measured LDOS becomes noticeably asymmetric with respect to the gap frequency. This skewness is fully consistent with our prediction that the spectrum consists of symmetric and antisymmetric Lorentzian components (Figs. 3g–i).